OpenStudy (anonymous):

Which of the following is true? A logarithmic function is not the inverse of an exponential function. The base of a logarithmic function can be a negative number. The domain of a logarithmic function f(x) = log2x is all real numbers greater than zero. The range of the logarithmic function f(x) = log2x is all real numbers less than zero

OpenStudy (anonymous):

Dang everybody came to the question like leeches hahaha.

OpenStudy (bittuaryan):

First only...A logarithmic function is not the inverse of an exponential function.

OpenStudy (anonymous):

Can ya help with a log problem?

OpenStudy (anonymous):

or actually, how to you input a lowered number, like an opposite ^3

OpenStudy (anonymous):

onto an online calculator

OpenStudy (anonymous):

Which of the following is the solution of log (x + 3)0.001 = -3 ?

OpenStudy (anonymous):

the (x+3) is lowered though not to the power of

OpenStudy (anonymous):

\[_{3}\] like that but i need it in a seperate calculator -______-

OpenStudy (anonymous):

what is it called?

OpenStudy (anonymous):

Which of the following is the solution of \[\log _{x+3}0.001 = -3 ?\]

OpenStudy (bittuaryan):

0.001 = 1/1000 = 10^-3 log[x + 3](0.001) = -3 log[x + 3](10^-3) = -3 REWRITE THIS IN EXPONENTIAL form 10^-3 = (X + 3)^-3 since we have the same exponents we can equate the base. 10 = x + 3 10 -3 = x x = 7

zepdrix (zepdrix):

@bittuaryan I'm confused what you were saying about the first question. Option A is clearly false. The question is asking for which one is true.

OpenStudy (anonymous):

ooo

zepdrix (zepdrix):

The logarithm and exponential ARE inverse functions of one another. :D The log function looks something like this. I think we're looking for option .... C probably. The domain represents legal \(x\) values the function can take on. Hmm it looks like this line is defined for all positive values of x yes? In other words the domain of a log function will be ~ All reals greater than 0.